For each system, choose the best description of its solution(no solution, unique solution, infinite solutions). If applicable, give the solution.
x+2y=2 And -x-2y=6
-x-2y=-2 x+2y=6

Found 2 solutions by Boreal, Alan3354:
Answer by Boreal(15235)

x+2y=2
-x-2y=6
0+0=8. No solution.
==========
-x-2y=-2
x+2y=6
0=4 no solution

Answer by Alan3354(69443)

For each system, choose the best description of its solution(no solution, unique solution, infinite solutions). If applicable, give the solution.
x+2y=2 And -x-2y=6
-x-2y=-2 x+2y=6
================
x+2y=2
-x-2y=-2
---------
The 2nd eqn is -1 times the 1st --> dependent --> infinite solutions.
========================================
-x-2y=6
x+2y=6
--------------- Add
0 = 12 --> inconsistent --> no solution
==================================
It's not clear how the equations are paired. The other tutor paired them differently.
Don't enter them like that.